OpenStudy (anonymous):

how to sketch y = 3 cos theta - 4

OpenStudy (dumbcow):

Amplitude = 3 period = 2pi vertical shift = -4 max = -4 +3 = -1 min = -4 -3 = -7 start at point (0,-1) next points at theta = pi/2 , pi, 3pi/2, 2pi

OpenStudy (dumbcow):

use a unit circle

OpenStudy (ddcamp):

It's a lot like cos(theta), just taller and shifted down 4. The y intercept would be -1, and the minimum would be -7. Minimums and maximums would happen at the same place as on cos(theta)